OpenStudy (angela210793):

I got 60 but then I saw how my teacher had solved it and she had got 75,I asked her today for the same q. and she got 57......What's the answer? Oo How many even,three digit numbers can be formed by using no.s from A={0;1;2;3;4;5} ****without repetition*****

OpenStudy (anonymous):

\[3\times 4P2?\]

OpenStudy (anonymous):

I would say the three digit numbers that can be formed are \(6(5)(4)=120\). The even numbers are then \(60\). Not entirely sure though.

OpenStudy (angela210793):

I got 5*4*3=60 It can be even only if the last digit is 0,2, or 4 so 3 choices then the 1st digit can't be 0 so 4 choices and 5 for the 2nd digit...tha'ts wht I thought....

OpenStudy (anonymous):

the last digit has to be a either a 0 or 2 or 4 for it to be an even number. so 3 ways you can choose the last digit. next you have 5 ways to choose the second last digit. next you have 4 ways of choosing the first digit. giving you 4*5*3=60

OpenStudy (anonymous):

Oh sorry, should be \(3\times 5P2\) which as you say, equals 60.

OpenStudy (angela210793):

Yahoo!!! :):):) I'm smarter than my teacher :P:P:P:P Thanks guys :):):)

OpenStudy (anonymous):

always challenge something if it doesnt make sense and you are welcome :)

OpenStudy (anonymous):

Probably best not to say that to her :P

OpenStudy (angela210793):

Hahha lol .....She can't do any harm to me now,I'm done with hs :P:P:P:P :)

OpenStudy (anonymous):

I wonder how she got 75.

OpenStudy (anonymous):

hold on a minute. 60 is incorrect.

OpenStudy (angela210793):

she did 5*5*3

OpenStudy (angela210793):

but today she did 5*4*1+4*4*1+4*4*1=57

OpenStudy (angela210793):

why is incorrect Dhatraditya?

OpenStudy (anonymous):

That's strange, if you could repeat numbers I guess it would be 6*6*3 but I don't know where 5*5*3 comes from.

OpenStudy (angela210793):

52 not not 57 sorry**

OpenStudy (anonymous):

52 is right

OpenStudy (anonymous):

Why?

OpenStudy (anonymous):

Is the number 012 a 3-digit number?

OpenStudy (angela210793):

OMG!!! Which 1's right?

OpenStudy (anonymous):

with 0 as the last digit, you can have 5 numbers to choose from for the middle digit and 4 numbers to choose from for the first digit. with 2 and 4 as the last digit, however, you have 5 numbers to choose from for the middle digit, but you dont have 0 as an option to be the first digit, because then it would be just a two digit number.

OpenStudy (angela210793):

no it's not that's why there r left 4 no.s 4 the 1st digit

OpenStudy (anonymous):

Lets consider the case where the ending digit is zero The no of cases are 5 x 4 x 1 = 20 Now lets consider the cases where the ending digit is 2 or 4 The no of cases are 4(excluding zero) x 4 x 2 = 32 Totals 32

OpenStudy (angela210793):

huh? 1st 75 then 57 then 60 now 32????????

OpenStudy (anonymous):

oh srry 52..i meant 20 + 32

OpenStudy (angela210793):

Probably 52 makes more sense....

OpenStudy (anonymous):

yes it is 52

OpenStudy (angela210793):

oh....so the final answer is 52....?

OpenStudy (anonymous):

indeed..u understood the method?

OpenStudy (angela210793):

Yea I did...thnx :):):)

OpenStudy (anonymous):

no problems

OpenStudy (anonymous):

him1618, why did you do 4x4x2 for the other two? the middle digit can still be 0

OpenStudy (angela210793):

let me draw smth plsss

OpenStudy (anonymous):

the middle 4 includes a ZERO see the middle digit has 5 choices including zero, but one choice has already been used as the first digit, so actually only 4 ways remain

OpenStudy (anonymous):

No repetition remember

OpenStudy (anonymous):

012345 are 6 digits, so with one digit used up, 5 digits remain for the middle digit, including 0

OpenStudy (anonymous):

but uve used one from 2 or 4 already to fill up the last digit havent u?

OpenStudy (anonymous):

what I mean is, suppose you used 2 for the last digit, you still have 0,1,3,4,5 to choose from. that is 5 digits.

OpenStudy (anonymous):

for the middle digit of course

OpenStudy (anonymous):

uve already used 1 3 4 or 5 in the first digit..so that leaves u 4 choices

OpenStudy (anonymous):

right right. I got it. :)

OpenStudy (anonymous):

thanks

OpenStudy (angela210793):

This is what Him meant i guess

OpenStudy (anonymous):

yes

OpenStudy (angela210793):

Thnx :):)

OpenStudy (anonymous):

no prob :))))))))))